# Membership Inference for Contrastive Pre-training Models with Text-only PII Queries

### 15 Mar 2026 v1

@@ Abstract
Contrastive pretraining models such as CLIP and CLAP underpin many vision-language and audio-language systems, yet their reliance on web-scale data raises growing concerns about memorizing Personally Identifiable Information (PII). Auditing such models via membership inference is challenging in practice: shadow-model MIAs are computationally prohibitive for large multimodal backbones, and existing multimodal attacks typically require querying the target with paired biometric inputs, thereby directly exposing sensitive biometric information to the target model. We propose Unimodal Membership Inference Detector (UMID), a text-only auditing framework that performs text-guided cross-modal latent inversion and extracts two complementary signals, similarity (alignment to the queried text) and variability (consistency across randomized inversions). UMID compares these statistics to a lightweight non-member reference constructed from synthetic gibberish and makes decisions via an ensemble of unsupervised anomaly detectors. Comprehensive experiments across diverse CLIP and CLAP architectures demonstrate that UMID significantly improves the effectiveness and efficiency over prior MIAs, delivering strong detection performance with sub-second auditing cost while complying with realistic privacy constraints.
keywords:
Membership Inference, Contrastive Pre-training, Privacy Leakage, Anomaly Detection
fn1fn1footnotetext: Co-first authors. Corresponding author. E-mail: huangyiyan@gbu.edu.cn
[…]
[besti]organization=Beijing Electronic Science and Technology Institute, city=Beijing, postcode=100070, country=China
[…]
[ali]organization=Alibaba Group, city=Beijing, postcode=100027, country=China
[…]
[ntu]organization=Nanyang Technological University, city=Singapore, postcode=637371, country=Singapore
[…]
[gbu]organization=School of Computing and Information Technology, Great Bay University, city=Dongguan, state=Guangdong, postcode=523000, country=China
1 Introduction
Contrastive pretraining models [53], exemplified by Contrastive Language–Image Pretraining (CLIP) [35] and Contrastive Language–Audio Pretraining (CLAP) [10], have become foundational backbones for modern vision–language and audio–language systems. Their strong performance, however, is largely enabled by web-scale training data that can contain Personally Identifiable Information (PII) [38, 41], raising increasing concerns about PII leakage and downstream misuse [50, 14]. Recent studies further suggest that multimodal models may inadvertently memorize membership signals or identity attributes, exposing individuals to privacy risks when sensitive biometric data (e.g., faces or voices) appear in the training set [54]. Consequently, auditing these multimodal models for membership leakage has become a critical imperative, not only to meet regulatory requirements, but also to provide an accessible safeguard that enables the public to assess potential data exposure in widely deployed foundation models without technical barriers [24].
@@ 1 Introduction
These challenges underscore an urgent need for auditing methods that avoid querying the target model with matched multimodal pairs, in line with the principle of multimodal data protection [27]. To this end, we study MIAs in a more realistic identity-level auditing setting. Specifically, we assume each training sample consists of a textual PII description $t_{i}$ (e.g., a name) and a corresponding non-textual modality $m_{i}$ (e.g., a face image for CLIP or a voice recording for CLAP). Given only a target text query $t$, the auditor seeks to determine whether there exists any training sample $(t_{i},m_{i})$ such that $t_{i}=t$. Crucially, this must be done under a strict unimodal constraint: even if the auditor possesses real biometric samples, they are prohibited from submitting them to the target model. This restriction enables lightweight, text-only queries that mitigate privacy leakage risk and are practical for third-party deployment, as shown in Figure 2.
@@
Figure 1: Comparison of UMID and traditional MIA methods. Traditional approaches rely on shadow models or bimodal input, while UMID turns membership inference problem into anomaly detection using text-only queries.
@@
Figure 2: Visualization of geometric separation. The extracted similarity and variability features via latent inversion exhibit a clear distributional gap between samples within and outside the training dataset of the target model.
@@
In this paper, we propose Unimodal Membership Inference Detector (UMID), a text-only query framework that detects PII memorization in contrastive models without exposing sensitive biometric inputs or training costly shadow models. Given a target text description, UMID performs cross-modal latent inversion with multiple randomized initializations and extracts two key statistics: (1) similarity, quantified as the average cosine alignment between the optimized embedding and the queried text embedding; and (2) variability, measured by the mean squared dispersion of the optimized embeddings across runs. UMID then compares these statistics against a lightweight non-member baseline built from synthetic textual “gibberish” and makes a membership decision using a voting ensemble of anomaly detectors. The intuition behind this design is simple yet heuristic: texts that correspond to identities seen during training (members) tend to produce more accurate and more consistent modal reconstructions under inversion than unseen texts (non-members), creating a clear distributional separation that can be detected reliably, as illustrated in Figure 2.
@@ 2 Related Work
The unprecedented capabilities of Multimodal Large Language Models (MLLMs) stem from extensive pretraining on massive and weakly supervised datasets [2]. Such model training process, however, inevitably introduces privacy risks as these web-scale datasets frequently contain personally identifiable information (PII), including faces, names, and voiceprints [30, 46]. A growing number of studies demonstrate that multimodal encoders and associated LLM components can memorize identity-linked features and sensitive attributes, rendering them susceptible to various privacy attacks such as model stealing, knowledge extraction, data reconstruction, and membership inference [22]. These risks are especially pronounced in contrastive frameworks such as CLIP, where strong cross-modal alignment allows adversaries to elicit sensitive identity associations through crafted queries [21]. Consequently, developing rigorous auditing protocols to detect such representation-level PII leakage remains a critical open problem [34].
@@
Membership inference attacks (MIAs) serve as a fundamental and widely-used instrument for quantifying privacy leakage by ascertaining whether a specific data sample was included in a model’s training set [40]. Conventional MIA pipelines typically rely on the training of shadow models to approximate the target model’s decision boundaries [40] or the exploitation of loss-based confidence signals [48]. However, as architectures of modern large models usually scale to billions of parameters, these strategies have become computationally prohibitive and often impractical [52]. Recent efforts have adapted MIAs to contrastive models such as CLIP by leveraging cross-modal alignment behavior, including identity detection inference attacks (IDIA) [13], cosine similarity attacks (CSA) [22], and weak supervision attacks (WSA) [36]. While these approaches successfully reduce or eliminate the need for shadow training, they paradoxically necessitate querying the target model with biometric inputs (e.g., facial images or voice recordings) during inference [43]. This highlights an urgent need for novel membership detectors that can operate effectively without involving sensitive bimodal PII.
3 Method
3.1 Problem Formulation
@@
Consider a multimodal contrastive pretraining model $\mathcal{M}$ (e.g., CLIP [20] or CLAP [49]) pre-trained on a dataset $D_{\text{train}}$. Each training sample $s_{i}=(t_{i},m_{i})$ represents a paired identity (i.e., PII) of an individual, where $t_{i}$ denotes a textual description (e.g., a person name) and $m_{i}$ denotes a corresponding non-textual modality (e.g., a face image for CLIP or a voice recording for CLAP). The objective of the detector is to perform identity-level membership inference: given a target identity $t$, the detector aims to determine whether the PII sample $(t,m)$ was present in $\mathcal{D}_{\text{train}}$. Notably, this is conducted under a unimodal privacy constraint: the auditor queries the model $\mathcal{M}$ using only textual descriptions. Even if the auditor possesses real biometric samples of the target, they are prohibited from submitting them to $\mathcal{M}$ to prevent secondary privacy leakage, which we term as the auditor’s dilemma in Section 1.
@@
We operate under a gray-box model auditing setting, where the detector is assumed to have query and gradient access to the frozen encoders, $\phi_{\text{text}}$ and $\phi_{\text{mod}}$ to perform the latent inversion described in Section 3.2. Crucially, this assumption does not require access to the original training pipeline: the auditor remains completely oblivious to the private training dataset $\mathcal{D}_{\text{train}}$, the specific training hyperparameters, or the optimizer states. This setting aligns with standard safety auditing protocols for open-weight foundation models, where model parameters are accessible but training data remains proprietary.
@@
3.2 UMID: Unimodal Membership Inference Detector
@@
    $D_{n}^{2}(t):=\frac{1}{n}\sum_{i=1}^{n}\|v^{(i)}(t)-\bar{v}_{n}(t)\|_{2}^{2},\quad\text{where }\bar{v}_{n}(t)=\frac{1}{n}\sum_{i=1}^{n}v^{(i)}(t).$ (2)
@@
Theorem 3.1 (Finite-sample geometric separation).
@@
$S_{n}(t_{\mathrm{in}})\geq s_{\mathrm{thr}},\ D_{n}^{2}(t_{\mathrm{in}})\leq d_{\mathrm{thr}}^{2},\qquad S_{n}(t_{\mathrm{out}})\leq s_{\mathrm{thr}},\ D_{n}^{2}(t_{\mathrm{out}})\geq d_{\mathrm{thr}}^{2}.$ (3)
Theorem 3.1 formally validates that our randomized reverse-optimization operationalizes the geometric distinction described above. It guarantees that the empirical statistics $(S_{n},D_{n}^{2})$ inherit the population-level separation with high probability. Furthermore, it quantifies the sample complexity, highlighting that the computational cost scales inversely with the square of the separation margin $\Gamma$. Detailed proofs are deferred to B.
@@
To approximate the non-member distribution, we generate a reference set of $\ell$ gibberish strings $\mathcal{G}=\{g_{1},\ldots,g_{\ell}\}$. Since $\mathcal{G}$ contains no semantic information and did not appear in training, it satisfies the isotropy condition stated in Section 3.2 and serves as a typical proxy for the null hypothesis.
@@ 4.1 Experimental Setup
We utilize LibriSpeech [33] and construct a richer speaker recognition dataset based on CommonVoice18.0 [1]. The latter involves 3,000 speakers (1,500 members/1,500 non-members) with detailed PII descriptions (ID, age, gender) augmented by GPT-4o to generate semantic background narratives. Similarly, we evaluate under One-Shot (1 audio/user) and Many-Shot (50 audios/user) settings.
@@
For CLIP, we conduct membership inference on six target models spanning three visual backbones: ResNet-50, ResNet-50x4, and ViT-B/32. The ResNet-based models follow the standard ResNet architecture [12, 45], while ViT-B/32 adopts the Vision Transformer design, covering both CNN- and Transformer-based encoders. For CLAP, we adopt a dual-encoder architecture in which the audio encoder is HTSAT [6], a transformer composed of four groups of Swin Transformer blocks [29], and the text encoder is RoBERTa [28]. The penultimate-layer outputs of both encoders (768 dimensions) are projected to a shared 512-dimensional embedding space via a two-layer MLP with ReLU activation. All target models are trained from scratch on their respective member datasets to establish unambiguous ground-truth membership labels. When real facial images are available for optional enhancement, DeepFace [39] is used as an external feature extractor.
@@
Implementation and Evaluation
Table 1: Comparison of CLIP (image-text) methods. Best and second-best results are bolded.
@@
4.2 Comparison with State-of-the-Art Methods
Performance on vision-language models
[…]
Performance on audio-language models
Table 2 compares CLAP membership auditing methods across two datasets, revealing three consistent findings: (i) Effectiveness and robustness: UMID forms the clear top tier in all settings, uniquely maintaining high precision and high recall simultaneously. On LibriSpeech, UMID already performs strongly without any real target audio (UMID-Null: $91.27\%/93.07\%$ accuracy in one-/many-shot), exceeding the strongest baseline AuditMI by a clear margin. Notably, with only one audio query, the gap further widens (up to +7.08+7.08 accuracy points) and recall approaches saturation. This advantage is amplified on the more challenging CommonVoice dataset, where all baselines degrade markedly (e.g., AuditMI $\approx 75\%$ accuracy) but UMID remains robust, widening both accuracy and recall gaps, indicating better generalization beyond clean speech. (ii) Exposure sensitivity. UMID’s one-shot performance is already close to many-shot (accuracy gaps $\leq 2.2$ points), suggesting that membership leakage is detectable even under minimal exposure when cross-modal evidence is properly exploited. (iii) Efficiency and deployability. In addition to its strong effectiveness, UMID runs in sub-second time ($\approx 0.6$$0.8$s), offering an order-of-magnitude speedup over other baselines.
Table 2: Comparison of CLAP (audio-text) methods. Best and second-best results are bolded.
@@ 4.2 Comparison with State-of-the-Art Methods
4.3 Ablation Study

### 22 Apr 2026 v2

@@ Abstract
Contrastive pretraining models such as CLIP and CLAP, serve as the ubiquitous perceptual backbones for modern multimodal large models, yet their reliance on web-scale data raises growing concerns about memorizing Personally Identifiable Information (PII). Auditing such models via membership inference is challenging in practice: shadow-model MIAs are computationally prohibitive for large multimodal backbones, and existing multimodal auditing methods typically require querying the target with paired biometric inputs, thereby directly exposing sensitive biometric information to the target model. To bypass this critical limitation, we demonstrate a highly desirable capability for privacy auditing: multimodal memorization within these foundational encoders can be accurately inferred using exclusively the text modality. We propose Unimodal Membership Inference Detector (UMID), a text-only auditing framework that performs text-guided cross-modal latent inversion and extracts two complementary signals, similarity (alignment to the queried text) and variability (consistency across randomized inversions). UMID compares these statistics to a lightweight non-member reference constructed from synthetic gibberish and makes decisions via an ensemble of unsupervised anomaly detectors. Comprehensive experiments across diverse CLIP and CLAP architectures demonstrate that UMID significantly improves the effectiveness and efficiency over prior MIAs, delivering strong detection performance with sub-second auditing cost using solely text queries, completely circumventing the need for biometric inputs and complying with strict privacy constraints.
I Introduction
Contrastive foundation models [63], exemplified by CLIP [44] and CLAP [17], have become the ubiquitous perceptual backbones for modern Multimodal Large Language Models (MLLMs) [12] and various vision/audio–language systems. Their strong performance, however, is largely enabled by web-scale training data that can contain Personally Identifiable Information (PII) [47, 50], raising increasing concerns about PII leakage and downstream misuse [59, 21, 42, 65]. Recent studies further suggest that multimodal models may inadvertently memorize membership signals or identity attributes, exposing individuals to privacy risks when sensitive biometric data (e.g., faces or voices) appear in the training set [64]. Consequently, auditing these multimodal models for membership leakage has become a critical imperative, not only to meet regulatory requirements, but also to provide an accessible safeguard that enables the public to assess potential data exposure in widely deployed foundation models without technical barriers [31].
@@ I Introduction
To resolve the auditor’s dilemma, a text-only, unimodal gray-box auditing setting is necessitated by two primary compliance scenarios: (1) Regulatory Auditing: Agencies auditing open-weight models for PII (e.g., under the EU AI Act) often have weight access but are legally prohibited from handling citizens’ raw biometric data. (2) Third-party Compliance: External auditors may receive model weights yet remain barred from sensitive bimodal user databases by regulations like GDPR/HIPAA. UMID addresses these by eliminating the need for both computationally expensive shadow models and privacy-violating bimodal queries.
@@
These challenges underscore an urgent need for auditing methods that avoid querying the target model with matched multimodal pairs [34]. In this paper, we reveal a key insight that resolves this dilemma: the presence of PII in the multimodal training set can be accurately audited using exclusively the text modality. By exploiting the cross-modal alignment behavior of these foundational encoders, we demonstrate that privacy-preserving, single-modality probing is sufficient to detect severe memorization. To this end, we study MIAs in a more realistic identity-level auditing setting. Specifically, we assume each training sample consists of a textual PII description $t_{i}$ (e.g., a name) and a corresponding non-textual modality $m_{i}$ (e.g., a face image for CLIP or a voice recording for CLAP). Given only a target text query $t$, the auditor seeks to determine whether there exists any training sample $(t_{i},m_{i})$ such that $t_{i}=t$. Crucially, this must be done under a strict unimodal constraint: even if the auditor possesses real biometric samples, they are prohibited from submitting them to the target model. This restriction enables lightweight, text-only queries that mitigate privacy leakage risk and are practical for third-party deployment (Figure 1a), proving that auditors can effectively assess multimodal privacy risks without ever exposing sensitive biometric data.
@@
(a) Comparison of UMID and traditional MIA methods. Traditional approaches rely on costly shadow models or bimodal inputs (risking secondary privacy leakage), whereas UMID safely formulates membership inference as anomaly detection using text-only queries.
@@
(b) Visualization of geometric separation. The extracted similarity and variability features via latent inversion exhibit a clear distributional gap between samples within and outside the training dataset of the target model.
Figure 1: Overview of the UMID auditing framework and the resulting distributional gap. The UMID method enables text-only membership inference (a) and achieves clear geometric separation between training and non-training samples (b).
@@
In this paper, we propose Unimodal Membership Inference Detector (UMID), a text-only query framework that detects PII memorization in contrastive models without exposing sensitive biometric inputs or training costly shadow models. Given a target text description, UMID performs cross-modal latent inversion with multiple randomized initializations and extracts two key statistics: (1) similarity, quantified as the average cosine alignment between the optimized embedding and the queried text embedding; and (2) variability, measured by the mean squared dispersion of the optimized embeddings across runs. UMID then compares these statistics against a lightweight non-member baseline built from synthetic textual “gibberish” and makes a membership decision using a voting ensemble of anomaly detectors. The intuition behind this design is straightforward yet effective: texts corresponding to identities seen during training (members) yield more accurate and consistent modal reconstructions under inversion than unseen texts (non-members). This discrepancy creates a distinct distributional separation that can be reliably detected, as illustrated in Figure 1b.
@@ II Related Work
The unprecedented capabilities of Multimodal Large Language Models (MLLMs) stem from extensive pretraining on massive and weakly supervised datasets [2, 10, 41, 60]. Such model training processes, however, inevitably introduces privacy risks as these web-scale datasets frequently contain personally identifiable information (PII), including faces, names, and voiceprints [37, 55]. A growing number of studies demonstrate that multimodal encoders and associated LLM components can memorize identity-linked features and sensitive attributes, rendering them susceptible to various privacy attacks such as model stealing, knowledge extraction, data reconstruction, and membership inference [29]. These risks are especially pronounced in contrastive frameworks such as CLIP, where strong cross-modal alignment allows adversaries to elicit sensitive identity associations through crafted queries [28]. Consequently, developing rigorous auditing protocols to detect such representation-level PII leakage remains a critical open problem [43].
@@
Membership inference attacks (MIAs) serve as a fundamental and widely-used instrument for quantifying privacy leakage by ascertaining whether a specific data sample was included in a model’s training set [49]. Conventional MIA pipelines typically rely on the training of shadow models to approximate the target model’s decision boundaries [49] or the exploitation of loss-based confidence signals [57]. However, as architectures of modern large models usually scale to billions of parameters [11, 16, 3, 13], these strategies have become computationally prohibitive and often impractical [62]. Recent efforts have adapted MIAs to contrastive models such as CLIP by leveraging cross-modal alignment behavior, including identity detection inference attacks (IDIA) [20], cosine similarity attacks (CSA) [29], and weak supervision attacks (WSA) [45]. While these approaches successfully reduce or eliminate the need for shadow training, they paradoxically necessitate querying the target model with biometric inputs (e.g., facial images or voice recordings) during inference [52]. This highlights an urgent need for novel membership detectors that can operate effectively without involving sensitive bimodal PII.
III Method
III-A Problem Formulation
@@
Consider a multimodal contrastive pretraining model $\mathcal{M}$ (e.g., CLIP [27] or CLAP [58]) pre-trained on a dataset $D_{\text{train}}$. Each training sample $s_{i}=(t_{i},m_{i})$ represents a paired identity (i.e., PII) of an individual, where $t_{i}$ denotes a textual description (e.g., a person name) and $m_{i}$ denotes a corresponding non-textual modality (e.g., a face image for CLIP or a voice recording for CLAP). The objective of the detector is to perform identity-level membership inference: given a target identity $t$, the detector aims to determine whether the PII sample $(t,m)$ was present in $\mathcal{D}_{\text{train}}$. Notably, this is conducted under a unimodal privacy constraint: the auditor queries the model $\mathcal{M}$ using only textual descriptions. Even if the auditor possesses real biometric samples of the target, they are prohibited from submitting them to $\mathcal{M}$ to prevent secondary privacy leakage, which we term as the auditor’s dilemma in Section I.
@@
We operate under a gray-box model auditing setting, where the detector is assumed to have query and gradient access to the frozen encoders, $\phi_{\text{text}}$ and $\phi_{\text{mod}}$ to perform the latent inversion described in Section III-B. Crucially, this assumption does not require access to the original training pipeline: the auditor remains completely oblivious to the private training dataset $\mathcal{D}_{\text{train}}$, the specific training hyperparameters, or the optimizer states. This setting aligns with standard safety auditing protocols for open-weight foundation models, where model parameters are accessible but training data remains proprietary.
@@
III-B UMID: Unimodal Membership Inference Detector
@@
    $\begin{split}D_{n}^{2}(t):=&\frac{1}{n}\sum_{i=1}^{n}\|v^{(i)}(t)-\bar{v}_{n}(t)\|_{2}^{2},\\ &\text{where }\bar{v}_{n}(t)=\frac{1}{n}\sum_{i=1}^{n}v^{(i)}(t).\end{split}$ (2)
@@
Theorem III.1 (Finite-sample geometric separation).
@@
$S_{n}(t_{\mathrm{in}})\geq s_{\mathrm{thr}},\quad D_{n}^{2}(t_{\mathrm{in}})\leq d_{\mathrm{thr}}^{2},$ (3)
$S_{n}(t_{\mathrm{out}})\leq s_{\mathrm{thr}},\quad D_{n}^{2}(t_{\mathrm{out}})\geq d_{\mathrm{thr}}^{2}.$
Theorem III.1 formally validates that our randomized reverse-optimization operationalizes the geometric distinction described above. It guarantees that the empirical statistics $(S_{n},D_{n}^{2})$ inherit the population-level separation with high probability. Furthermore, it quantifies the sample complexity, highlighting that the computational cost scales inversely with the square of the separation margin $\Gamma$. Detailed proofs are deferred to B.
@@
To approximate the non-member distribution, we generate a reference set of $\ell$ gibberish strings $\mathcal{G}=\{g_{1},\ldots,g_{\ell}\}$. Since $\mathcal{G}$ contains no semantic information and did not appear in training, it satisfies the isotropy condition stated in Section III-B and serves as a typical proxy for the null hypothesis.
@@ IV-A Experimental Setup
We utilize LibriSpeech [40] and construct a richer speaker recognition dataset based on CommonVoice18.0 [1]. The latter involves 3,000 speakers (1,500 members/1,500 non-members) with detailed PII descriptions (ID, age, gender) augmented by GPT-4o to generate semantic background narratives. Similarly, we evaluate under One-Shot (1 audio clip/user) and Many-Shot (50 audio clips/user) settings.
@@
For CLIP, we conduct membership inference on six target models spanning three visual backbones: ResNet-50, ResNet-50x4, and ViT-B/32. The ResNet-based models follow the standard ResNet architecture [19, 54], while ViT-B/32 adopts the Vision Transformer design, covering both CNN- and Transformer-based encoders. For CLAP, we adopt a dual-encoder architecture in which the audio encoder is HTSAT [8], a transformer composed of four groups of Swin Transformer blocks [36], and the text encoder is RoBERTa [35]. The penultimate-layer outputs of both encoders (768 dimensions) are projected to a shared 512-dimensional embedding space via a two-layer MLP with ReLU activation. All target models are trained from scratch on their respective member datasets to establish unambiguous ground-truth membership labels. When real facial images are available for optional enhancement, DeepFace [48] is used as an external feature extractor. Following standard MIA protocols [5], we train models from scratch to ensure absolute ground-truth certainty. Auditing public foundation models (e.g., LAION-trained CLIP) is methodologically unreliable due to massive data contamination and unindexed training sets, which obscure membership labels. Our controlled training on industry-standard architectures (ResNet, ViT, RoBERTa) ensures that evaluation metrics reflect true algorithmic capability rather than dataset noise.
@@
TABLE I: Comparison of CLIP (image-text) methods. Best and second-best results are bolded.
@@
TABLE II: Comparison of CLAP (audio-text) methods. Best and second-best results are bolded.
@@
Implementation and Evaluation
IV-B Comparison with State-of-the-Art Methods
Performance on vision-language models
[…]
Performance on audio-language models
Table II compares CLAP membership auditing methods across two datasets, revealing three consistent findings: (i) Effectiveness and robustness: UMID forms the clear top tier in all settings, uniquely maintaining high precision and high recall simultaneously. On LibriSpeech, UMID already performs strongly without any real target audio (UMID-Null: $91.27\%/93.07\%$ accuracy in one-/many-shot), exceeding the strongest baseline AuditMI by a clear margin. Notably, with only one audio query, the gap further widens (up to +7.08+7.08 accuracy points) and recall approaches saturation. This advantage is amplified on the more challenging CommonVoice dataset, where all baselines degrade markedly (e.g., AuditMI $\approx 75\%$ accuracy) but UMID remains robust, widening both accuracy and recall gaps, indicating better generalization beyond clean speech. (ii) Exposure sensitivity. UMID’s one-shot performance is already close to many-shot (accuracy gaps $\leq 2.2$ points), suggesting that membership leakage is detectable even under minimal exposure when cross-modal evidence is properly exploited. (iii) Efficiency and deployability. In addition to its strong effectiveness, UMID runs in sub-second time ($\approx 0.6$$0.8$s), offering an order-of-magnitude speedup over other baselines.